For a geometric sequence, t3 = 6 and ts = 2/81. Find t6



1
Expert's answer
2021-02-18T13:51:26-0500
"t_3=\dfrac{a}{b}"

"t_8=ar^7=\dfrac{2}{81}"

"\dfrac{t_8}{t_3}=\dfrac{ar^7}{ar^2}=r^5=\dfrac{\dfrac{2}{81}}{6}=\dfrac{1}{3^5 }"

"r=\dfrac{1}{3}"

"a(\dfrac{1}{3})^2=6=>a=54"

"t_6=54(\dfrac{1}{3})^5=\dfrac{2}{9}"